OpenStudy (anonymous):

\(*=^=*)/ please help .Use the quadratic formula to solve the quadratic equation x2 − 8x − 4 = 0.

OpenStudy (anonymous):

give me a sec.

zepdrix (zepdrix):

Hi there c: \(\Large\rm ax^2+bx+c=0\) Notice that our quadratic formula has all pluses, no minus. So we'll want to write our quadratic like this: \(\Large\rm x^2+(-8)x+(-4)\) so we can more easily identify our b and c.

zepdrix (zepdrix):

\(\Large\rm x=\dfrac{-(-8)\pm\sqrt{(-8)^2-4(1)(-4)}}{2(1)}\)Understand how I plugged all those values into the formula?

OpenStudy (anonymous):

thanks\ (*>v<)/

zepdrix (zepdrix):

Your emotes are kinda scary :o The first one looks like a creepy spider lolol

OpenStudy (anonymous):

lol